Select the ordered pair from the choices below that is a solution to the following system of equations:
4y = 2x + 10
8x - 3y = -14

Answer by benazir.sj@gmail.com(109)

4y = 2x + 10
2x-4y= -10---------------(1)
8x - 3y = -14--------------(2)
(1)*4 =4*2x -4*4 =-4*10
8x-16y=-40 (1a)
equate (1a) and (2)
(1a) - (2)
8x-16y = -40
8x-3y = -14
=> -13y= -26
y=-26/-13
y=2
substitute the value of y in(1)
4y=2x+10
4*2=2x+10
8=2x+10
8-10=2x
-2=2x
x=-2/2=-1
there fore the ordered pairs are
(x,y) =(-1,2)